The radius of a can soup is 2 in, while the height is 7 in.
What is the amount of soup this can holds? (Use pi = 22/7)
Provide a mathematical explanation of your answer.

Answer by RAY100(1637)

Volume = Area of Base * Height = pi * r^2 * hg = 22/7 * 2^2 *7 = 88